# Power Variable Projection for Initialization-Free Large-Scale Bundle Adjustment

### v1

@@ Abstract
Initialization-free bundle adjustment (BA) remains largely uncharted. While Levenberg-Marquardt algorithm is the golden method to solve the BA problem, it generally relies on a good initialization. In contrast, the under-explored Variable Projection algorithm (VarPro) exhibits a wide convergence basin even without initialization. Coupled with object space error formulation, recent works have shown its ability to solve (small-scale) initialization-free bundle adjustment problem. We introduce Power Variable Projection (PoVar), extending a recent inverse expansion method based on power series. Importantly, we link the power series expansion to Riemannian manifold optimization. This projective framework is crucial to solve large-scale bundle adjustment problem without initialization. Using the real-world BAL dataset, we experimentally demonstrate that our solver achieves state-of-the-art results in terms of speed and accuracy. In particular, our work is the first, to our knowledge, that addresses the scalability of BA without initialization and opens new venues for initialization-free Structure-from-Motion.
@@ 1 Introduction
Bundle adjustment (BA) is the key component of many structure-from-motion and 3D reconstruction algorithms. With the recent emergence of large-scale internet photo collections [3] and new applications (mixed reality, autonomous driving), the need to solve large-scale BA has become a hard challenge. Traditional BA formulation addresses the following question: given image measurements and approximate landmark positions and camera parameters, can we derive the exact positions and parameters? The gold standard is to use the Levenberg-Marquardt algorithm [27] coupled with the Schur complement trick and a scalable solver for the reduced camera system, which is often the preconditioned conjugate gradient algorithm. Latest work [25] achieves outstanding speed for large-scale BA by using a power series expansion of the inverse Schur complement.
Recently, a new line of works [13, 14, 11] has attempted to solve the BA problem without careful initialization: given only image measurements, how do we derive pose parameters and 3D landmark positions? This challenge is largely uncharted, and the scalability a blind spot. In particular, most of these works aim to formalize the problem into a stratified BA formulation, and none of them try to effectively design suitable solvers. It is noteworthy that even the most recent works only use direct factorization which becomes impractical for large-scale problems with several hundreds of cameras. The deficiency of competitive solvers, contrary to the traditional BA problem, can be broadly explained by the difference of convergence behaviours between a well-initialized problem and an initialization-free problem.
Following up on the recent findings concerning inverse expansion methods, we address the scalability of initialization-free BA. Our new solver based on the Variable Projection algorithm overcomes the issues of convergence of the scalable preconditioned conjugate gradients algorithm, while being very efficient for thousand of camera viewpoints. This leads to the following contributions:
@@
Figure 1: In contrast to traditional bundle adjustment problem, initialization-free BA is a very under-explored problem. It does not assume any approximation of pose and landmark parameters, making the problem much harder to solve. From a random initialization (left figure), and given only image measurements, we aim to recover pose and landmark parameters. Our approach, that extends inverse expansion method, is motivated by the lack of scalability of existing solvers. On the real-world BAL problems (e.g. Venice-89, right figure), we demonstrate the efficiency of the proposed combination of our novel solver Power Variable Projection (PoVar) and Riemannian manifold optimization framework for expansion method to solve the stratified BA problem.
@@ 2 Related Work
VarPro is an optimization approach for solving bivariate problems that can be formulated as minimizing a cost function $f(u,v)$ with $f:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ over two sets of variables $u\in\mathbb{R}^{m}$ and $v\in\mathbb{R}^{n}$. Unlike alternation, which fixes $u$ and optimizes over $v$ and vice versa, or joint optimization, which optimizes the stack of $u$ and $v$ simultaneously, variable projection replaces $v$ with $v^{*}(u):=\operatorname*{arg\,min}_{u}f(u,v)$ ($v$ always optimal over $u$) such that the optimized cost function $f(u,v^{*}(u))=:f^{*}(u)$ becomes a function of $u$ only. The original VarPro algorithm by Golub and Pereyra [9] and its approximations by Ruhe and Wedin [20] assume solving a separable nonlinear least squares (SNLS) problem where $v^{*}(u)$ can be obtained in closed form. VarPro was consistently ignored in the computer vision community and even misidentified as a form of alternating optimization [5]. It first receives proper attention when Okatani et al. [17] demonstrates VarPro equipped with a trust-region approach such as Levenberg-Marquardt can yield a wide basin of convergence for several toy problems that can be formulated as a SNLS problem such as affine structure-from-motion and factorization-based non-rigid structure-from-motion. Shortly after, Strelow [21, 22] extends VarPro to the nonlinear case where $v^{*}(u)$ is not in closed form such as bundle adjustment. Nevertheless, these works do not address the issue of increased algorithmic complexity. Later, it has been shown by Hong et al. [12] that VarPro can be efficiently implemented by performing inner iterations [2] (also known as embedded point iterations [15]) over the set of eliminated variables $u$ followed by performing a joint optimization step with no damping on $u$. While this enables faster runtime compared to previous studies, it has only been tested up to small-medium sized problems with around 300 camera views. To this date, no work to the best of our knowledge has improved the scalability of the VarPro algorithm beyond [12].
@@
While traditional (large-scale) bundle adjustment is regularly studied ([3, 6, 26, 7, 19, 30, 4]), initialization-free BA is a recent research topic. In a seminal work, Hong et al. [13] propose to solve projective bundle adjustment from arbitrary initialization with the Variable Projection algorithm. Notably, they propose a stratified bundle adjustment formulation with increasing difficulty. In a follow-up work, pOSE [11] incorporates the equivalence between nonlinear VarPro and the Schur complement recently identified in [12]. Additionally, they define an objective in-between affine and projective models that leads to a wide convergence bassin. Iglesias et al. [14] complements this pseudo object space error formulation with an exponential regularization term, and show interesting results on very-small-scale problems. Nevertheless, up to our knowledge, none of these studies look closely into the scalability of the proposed frameworks for initialization-free bundle adjustment.
@@
Despite its short recent appearance in the literature, inverse expansion method is a highly efficient and competitive approach for solving the linearized system of equations for bundle adjustment that is already challenging established factorization and iterative methods. It links the Schur complement [28] to the power series expansion of its inverse. Using the power Schur complement as a preconditioner for normal equations leads to very good results for physics problems such that convection-diffusion [29]. PoBA [25] proposes to directly apply the power Schur complement to the right-hand side of the reduced camera system. Their solver results in improved speed and accuracy for traditional bundle adjustment formulation with respect to the existing methods.
@@ 3 Problem Statement and Motivation
Recently, Equation 1 with good initialization has been solved very efficiently with a novel inverse expansion method.
@@ 3.1 Inverse expansion method
Inverse expansion method [29, 25] relies on the expansion of the inverse of a matrix into a power series, as stated in the following proposition:
@@
Let $M$ be a $n\times n$ matrix. If the spectral radius of $M$ satisfies $\lVert M\rVert<1$, then
@@
Given a linear approximation of Equation 1 followed by the Schur complement trick, Weber et al. [25] relies the inverse Schur complement of the Levenberg-Marquardt algorithm to its power series. They show significant improvement in terms of speed and accuracy to solve BA problem with good initialization. In addition to this empirical insights, some convergence behaviour concerning the approximated results for the BA problem are theoretically proved.
However, it is well-known in the literature (see e.g. [13]) that solving Equation 1 from arbitrary initialization is non-feasible. Hong et al. [11] override this challenge by proposing a stratified bundle adjustment problem. Let us revisit the formulation of this approach.
@@ 3.2 Initialization-free bundle adjustment
Given $n_{p}$ poses and $n_{l}$ landmarks, $x=(x_{p},x_{l})$ contains all the optimization variables. We mimic the camera as a projective model. For pose $i$, we consider the camera parameters $x_{p}^{i}\in\mathbb{R}^{3\times 4}$ and solve the following generic nonlinear separable problem:
$\min_{x_{p},\tilde{x}_{l}}F(x_{p},\tilde{x}_{l})=\big{\lVert}r(x_{p},\tilde{x}_{l})\big{\rVert}_{2}^{2}=\big{\lVert}G(x_{p})\tilde{x}_{l}^{j}-z(x_{p})\big{\rVert}_{2}^{2}\,,$ (5)
@@
with $x_{p}^{i,1:2}$ and $x_{p}^{i,3}$ respectively the first two lines and the third line of $x_{p}^{i}$, and $\eta\in[0,1]$.
@@
The cameras and landmarks parameters obtained by solving Equation 5 are refined by minimizing the projective standard objective [13] over the projective camera models in homogeneous form ($\{\tilde{x}_{p}^{i}~{}|~{}\mathrm{vec}(\tilde{x}_{p}^{i})\in\,S^{12}\}$) and the 3D landmarks in homogeneous coordinates ($\{\tilde{x}_{l}^{j}~{}|~{}\tilde{x}_{l}^{j}\in\,S^{4}\})$:
@@
The last stage is a minimization problem to enforce the projective camera matrices to satisfy $SE(3)$ properties. Note that our work mostly focuses on the first two stages. The proposed implementation for this third stage has illustrative purpose, see Figure 1. We refer the reader to Supplemental for further details.
@@ 3.3 Limitations and proposed method
In contrast to the Levenberg-Marquardt algorithm, very few solvers have been designed to efficiently solve VarPro. In practice, even the most recent works [11, 14] use a direct factorization (e.g. Cholesky decomposition, QR factorization), that is well-known to be very poorly scalable (see e.g. [3]), to solve Equation 5. In particular, we note that these works consider bundle adjustment problems with only few tens of cameras – sometimes less than ten. On the other hand, Hong and Fitzgibbon [10] show that coupling VarPro with the popular preconditioned conjugate gradients algorithm may not converge very efficiently.
We propose to build on recent inverse expansion method. We first adapt the power series expansion to the VarPro algorithm (Section 4.2). We show that the so-called PoVar efficiently solves Equation 5. Moving forward, we extend the power series expansion to Riemannian manifold optimization (Section 4.3). This new Riemannian framework, that we call RiPoBA, is necessary to use expansion method for solving Equation 7. We demonstrate that the combination of this two solvers is highly competitive (Section 5).
@@ 4 Power Variable Projection
First, let us briefly revisit the VarPro algorithm. We refer the reader to [9] for further details.
@@ 4.1 Revisited variable projection
In contrast to joint optimization, VarPro optimizes over landmark parameters and camera parameters, but in a way different from the standard alternating least squares such that landmarks are not assumed to be fixed when updating the camera parameters. It first considers Equation 5 as a nonlinear minimization problem over $x_{l}$ only. Due to the separability of the equation, a closed-form solution for optimal $x_{l}^{*}(x_{p})$ is straightforward:
$x_{l}^{*}(x_{p})=\operatorname*{arg\,min}_{\tilde{x}_{l}}\lVert G(x_{p})\tilde{x}_{l}-z(x_{p})\rVert_{2}^{2}=G(x_{p})^{{\dagger}}z(x_{p})\,,$ (8)
with $G(x_{p})^{{\dagger}}$ the pseudo-inverse of $G(x_{p})$. Substituting $x_{l}^{*}(x_{p})$ in Equation 5 leads to the following reduced problem:
@@
Following [12], Equation 9 can be solved with LM algorithm over $x_{p}$. The Jacobian of $r^{*}$ is approximated with the so-called RW2 approximation [16], that leads to the normal equation:
@@
with $J_{l}$ and $J_{p}$ respectively the landmark and pose Jacobians of the original residual (Equation 5) around an equilibrium $r^{0}$, and $D_{p}$ diagonal damping matrix for pose variables. In contrast to joint optimization, only the pose Jacobian is damped in the associated Hessian. It follows that $U_{\lambda}$ is symmetric positive-definite [23], whereas $V_{0}$ is only guaranteed to be symmetric positive-semidefinite. Nevertheless, we observe $V_{0}$ is usually of full rank unless a 3D landmark is observed by few cameras with narrow baselines. By using the Schur complement trick, the update equation for VarPro becomes:
@@
while sharing a close structure to the Schur complement of the traditional BA problem, has a very different convergence behaviour, due to the undamped landmark Jacobian $V_{0}$.
@@ 4.2 Power series for VarPro
Let $\mu$ be an eigenvalue of $U_{\lambda}^{-1}WV_{0}^{-1}W^{\top}$. Then
$\mu\in[0,1[\,.$ (16)
@@
It follows that the pose update $\Delta x_{p}$ in Equation 14 can be directly approximated by $x(m)$ with Proposition 1 applied to the Schur complement $S^{V}$:
@@
Once the pose update is estimated, the landmark update can be derived following closed-form Equation 8 or the incremental method described in [12]. That extends the inverse expansion method to Variable Projection algorithm. The difference between both solvers is the role of the damped parameters. While it seems slight, this is enough to offer a very dissimilar convergence behaviour, as we will see in the experiments.
Before that, let investigate Riemannian manifold optimization, that is necessary to solve the second stage of the stratified BA problem. In particular, we show that we can link this framework to expansion method.

### v5

# latexSupplementary-Literature

@@ Abstract
Most Bundle Adjustment (BA) solvers like the Levenberg-Marquardt algorithm require a good initialization. Instead, initialization-free BA remains a largely uncharted territory. The under-explored Variable Projection algorithm (VarPro) exhibits a wide convergence basin even without initialization. Coupled with object space error formulation, recent works have shown its ability to solve small-scale initialization-free bundle adjustment problem. To make such initialization-free BA approaches scalable, we introduce Power Variable Projection (PoVar), extending a recent inverse expansion method based on power series. Importantly, we link the power series expansion to Riemannian manifold optimization. This projective framework is crucial to solve large-scale bundle adjustment problems without initialization. Using the real-world BAL dataset, we experimentally demonstrate that our solver achieves state-of-the-art results in terms of speed and accuracy. To our knowledge, this work is the first to address the scalability of BA without initialization opening new venues for initialization-free structure-from-motion.
@@ 1 Introduction
Bundle adjustment (BA) is the key component of many structure-from-motion and 3D reconstruction algorithms. With the recent emergence of large-scale internet photo collections [3] and new applications (mixed reality, autonomous driving, digital twins), the need to solve large-scale BA has become an important challenge. Traditional BA addresses the following question: Given image measurements and approximate landmark positions and camera parameters, can we derive the exact positions and parameters? The gold standard is to use the Levenberg-Marquardt algorithm [25] coupled with the Schur complement trick and a scalable solver for the reduced camera system, which is often the preconditioned conjugate gradient algorithm. Recent work [23] achieves outstanding speed for large-scale BA by using a power series expansion of the inverse Schur complement.
Recently, a new line of works [13, 14, 11] has attempted to solve the BA problem without careful initialization: Given only image measurements, how do we derive pose parameters and 3D landmark positions? This challenge is largely uncharted, and its scalability a blind spot. In particular, most existing works aim to formalize the problem into a stratified BA formulation, and none of them try to design effective solvers. It is noteworthy that even the most recent works only use direct factorization which becomes impractical for large-scale problems with several hundreds of cameras. In contrast to the traditional BA problem, the deficiency of competitive solvers for initialization-free BA can be broadly explained by the difference of convergence behaviour between a well-initialized problem and an initialization-free problem.
Following up on the recent findings concerning inverse expansion methods, we address the scalability of initialization-free BA. Our new solver based on the Variable Projection algorithm overcomes the issues of convergence of the scalable preconditioned conjugate gradients algorithm, while being efficient for thousand of camera viewpoints. In summary, we make the following contributions:
@@
Figure 1: In contrast to traditional bundle adjustment problem, initialization-free BA is a largely under-explored problem. It does not assume any approximation of pose and landmark parameters, making the problem much harder to solve. From a random initialization (left figure), and given only image measurements, we aim to recover pose and landmark parameters. Our approach, that extends inverse expansion method, is motivated by the lack of scalability of existing solvers. On the real-world BAL problems (e.g. Venice-89, right figure), we demonstrate the efficiency of the proposed combination of our novel solver Power Variable Projection (PoVar) and Riemannian manifold optimization framework for expansion method to solve the stratified BA problem.
@@ 2 Related Work
VarPro is an optimization approach for solving bivariate problems that can be formulated as minimizing a cost function $f(u,v)$ with $f:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ over two sets of variables $u\in\mathbb{R}^{m}$ and $v\in\mathbb{R}^{n}$. Unlike alternation, which fixes $u$ and optimizes over $v$ and vice versa, or joint optimization, which optimizes the stack of $u$ and $v$ simultaneously, variable projection replaces $v$ with $v^{*}(u):=\operatorname*{arg\,min}_{v}f(u,v)$ ($v$ always optimal over $u$) such that the optimized cost function $f(u,v^{*}(u))=:f^{*}(u)$ becomes a function of $u$ only. The original VarPro algorithm by Golub and Pereyra [9] and its approximations by Ruhe and Wedin [19] assume solving a separable nonlinear least squares (SNLS) problem where $v^{*}(u)$ can be obtained in closed form. VarPro was consistently ignored in the computer vision community and even misidentified as a form of alternating optimization [5]. It first receives proper attention when Okatani et al. [17] demonstrates VarPro equipped with a trust-region approach such as Levenberg-Marquardt can yield a wide basin of convergence for several toy problems that can be formulated as a SNLS problem such as affine structure-from-motion and factorization-based non-rigid structure-from-motion. Shortly after, Strelow [20, 21] extends VarPro to the nonlinear case where $v^{*}(u)$ is not in closed form such as bundle adjustment. Nevertheless, these works do not address the issue of increased algorithmic complexity. Later, it has been shown by Hong et al. [12] that VarPro can be efficiently implemented by performing inner iterations [2] (also known as embedded point iterations [15]) over the set of eliminated variables $u$ followed by performing a joint optimization step with no damping on $u$. While this enables faster runtime compared to previous studies, it has only been tested up to small-medium sized problems with around 300 camera views. To this date, no work to the best of our knowledge has improved the scalability of the VarPro algorithm beyond [12].
@@
While traditional (large-scale) bundle adjustment is regularly studied ([3, 6, 24, 7, 18, 28, 4]), initialization-free BA is a recent research topic. In a seminal work, Hong et al. [13] propose to solve projective bundle adjustment from arbitrary initialization with the Variable Projection algorithm. Notably, they propose a stratified bundle adjustment formulation with increasing difficulty. In a follow-up work, pOSE [11] incorporates the equivalence between nonlinear VarPro and the Schur complement recently identified in [12]. Additionally, they define an objective in-between affine and projective models that leads to a wide convergence basin. Iglesias et al. [14] complements this pseudo object space error formulation with an exponential regularization term, and show interesting results on very-small-scale problems. Nevertheless, to our knowledge, none of these studies look closely into the scalability of the proposed frameworks for initialization-free bundle adjustment.
@@
Despite its short recent appearance in the literature, inverse expansion method is a highly efficient and competitive approach for solving the linearized system of equations for bundle adjustment that is already challenging established factorization and iterative methods. It links the Schur complement [26] to the power series expansion of its inverse. Using the power Schur complement as a preconditioner for normal equations leads to good results for physics problems such as convection-diffusion [27]. PoBA [23] proposes to directly apply the power Schur complement to the right-hand side of the reduced camera system. Their solver results in improved speed and accuracy for traditional bundle adjustment formulation with respect to the existing methods.
@@ 3 Problem Statement and Motivation
Recently, Eq. 1 with good initialization has been solved efficiently with a novel inverse expansion method.
@@ 3.1 Inverse expansion method
Inverse expansion method [27, 23] relates on the expansion of the inverse of a matrix into a power series, as stated in the following proposition:
@@
Let $M$ be an $n\times n$ matrix. If the spectral radius of $M$ satisfies $\lVert M\rVert<1$, then
@@
Given a linear approximation of Eq. 1 followed by the Schur complement trick, Weber et al. [23] relates the inverse Schur complement of the Levenberg-Marquardt algorithm to its power series. They show significant improvement in terms of speed and accuracy to solve BA problem with good initialization. In addition to this empirical insights, some convergence behaviour concerning the approximated results for the BA problem are theoretically proved.
However, it is well-known in the literature (see e.g. [13]) that solving Eq. 1 from arbitrary initialization is non-feasible. Hong et al. [11] override this challenge by proposing a stratified bundle adjustment problem. Let us revisit the formulation of this approach.
@@ 3.2 Initialization-free bundle adjustment
Given $n_{p}$ poses and $n_{l}$ landmarks, $x=(x_{p},x_{l})$ contains all the optimization variables. We model the camera as a projective one. For pose $i$, we consider the camera parameters $x_{p}^{i}\in\mathbb{R}^{3\times 4}$ and solve the following generic nonlinear separable problem:
$\min_{x_{p},\tilde{x}_{l}}F(x_{p},\tilde{x}_{l})=\big{\lVert}r(x_{p},\tilde{x}_{l})\big{\rVert}_{2}^{2}=\big{\lVert}G(x_{p})\tilde{x}_{l}-z(x_{p})\big{\rVert}_{2}^{2}\,,$ (5)
@@
with $x_{p}^{i,1:2}$ and $x_{p}^{i,3}$ respectively the first two rows and the third row of $x_{p}^{i}$, and $\eta\in[0,1]$.
@@
The cameras and landmarks parameters obtained by solving Eq. 5 are refined by minimizing the projective standard objective [13] over the projective camera models in homogeneous form ($\{\tilde{x}_{p}^{i}~{}|~{}\mathrm{vec}(\tilde{x}_{p}^{i})\in\,S^{12}\}$) and the 3D landmarks in homogeneous coordinates ($\{\tilde{x}_{l}^{j}~{}|~{}\tilde{x}_{l}^{j}\in\,S^{4}\})$:
@@
The last stage is a minimization problem to enforce the projective camera matrices to satisfy $SE(3)$ properties. Note that our work mostly focuses on the first two stages. The proposed implementation for this third stage has illustrative purpose, see Fig. 1. We refer the reader to Supplemental for further details.
@@ 3.3 Limitations and proposed method
In contrast to the Levenberg-Marquardt algorithm, few solvers have been designed to efficiently solve VarPro. In practice, even the most recent works [11, 14] use a direct factorization (e.g. Cholesky decomposition, QR factorization), that is well-known to be poorly scalable (see e.g. [3]), to solve Eq. 5. In particular, we note that these works consider bundle adjustment problems with only few tens of cameras – sometimes less than ten. On the other hand, Hong and Fitzgibbon [10] show that coupling VarPro with the popular preconditioned conjugate gradients algorithm may not converge efficiently.
We propose to build on recent inverse expansion method. We first adapt the power series expansion to the VarPro algorithm (Section 4.2). We show that the so-called PoVar efficiently solves Eq. 5. Moving forward, we extend the power series expansion to Riemannian manifold optimization (Section 4.3). This new Riemannian framework, that we call RiPoBA, is necessary to use expansion method for solving Eq. 7. We demonstrate that the combination of this two solvers is highly competitive (Section 5).
@@ 4 Power Variable Projection
We start by revisiting the VarPro algorithm. We refer to [9] for further details.
@@ 4.1 Revisited variable projection
In contrast to joint optimization, VarPro optimizes over landmark parameters and camera parameters, but in a way different from the standard alternating least squares such that landmarks are not assumed to be fixed when updating the camera parameters. It first considers Eq. 5 as a nonlinear minimization problem over $\tilde{x}_{l}$ only. Due to the separability of the equation, a closed-form solution for optimal $\tilde{x}_{l}^{*}(x_{p})$ is straightforward:
$\tilde{x}_{l}^{*}(x_{p})=\operatorname*{arg\,min}_{\tilde{x}_{l}}\lVert G(x_{p})\tilde{x}_{l}-z(x_{p})\rVert_{2}^{2}=G(x_{p})^{{\dagger}}z(x_{p})\,,$ (8)
with $G(x_{p})^{{\dagger}}$ the pseudo-inverse of $G(x_{p})$. As we set the last coordinate of $\tilde{x}_{l}$ to $1$, we normalize $\tilde{x}_{l}^{*}(x_{p})$ by its last coordinate and its substitution in Eq. 5 leads to the following reduced problem:
@@
Following [12], Eq. 9 can be solved with LM algorithm over $x_{p}$. The Jacobian of $r^{*}$ is approximated with the so-called RW2 approximation [16], that leads to the normal equation:
@@
with $J_{l}$ and $J_{p}$ respectively the landmark and pose Jacobians of the original residual (Eq. 5) around an equilibrium $r^{0}$, and $D_{p}$ diagonal damping matrix for pose variables. In contrast to joint optimization, only the pose Jacobian is damped in the associated Hessian. It follows that $U_{\lambda}$ is symmetric positive-definite [22], whereas $V_{0}$ is only guaranteed to be symmetric positive-semidefinite. Nevertheless, we observe $V_{0}$ is usually of full rank unless a 3D landmark is observed by few cameras with narrow baselines. By using the Schur complement trick, the update equation for VarPro becomes:
@@
while sharing a close structure to the Schur complement of the traditional BA problem, has a different convergence behaviour, due to the undamped landmark Jacobian $V_{0}$.
@@ 4.2 Power series for VarPro
Let $\mu$ be an eigenvalue of $U_{\lambda}^{-1}WV_{0}^{{\dagger}}W^{\top}$. Then
$0\leq\mu<1\,.$ (16)
@@
It follows that the pose update $\Delta x_{p}$ in Eq. 14 can be directly approximated [^1] As in practice $V_{0}$ is full-rank, and to avoid encumbering notations, we write $V_{0}^{-1}$ instead of the pseudo-inverse in the rest of the paper. by $x(m)$ with Proposition 1 applied to the Schur complement $S^{V}$:
@@
Once the pose update is estimated, the landmark update is derived following closed-form Eq. 8. That extends the inverse expansion method to Variable Projection algorithm. The difference between both solvers is the role of the damped parameters. While it seems slight, this is enough to offer a dissimilar convergence behaviour, as we will see in the experiments.
Before that, let us investigate Riemannian manifold optimization, that is necessary to solve the second stage of the stratified BA problem. In particular, we show that we can link this framework to expansion method.